OpenStudy (anonymous):

easy question easy medal x^2-16= (x+4)(x-4) Why can't it be (x+8)(x-2)?

OpenStudy (anonymous):

x+8 and x-2 evaluate and multiply not to give 2 terms, but three terms, respectively x^2 + 6x -16, instead of x^2-16.

OpenStudy (anonymous):

oh so how am i supposed to know that during these problems is there a specific method?

OpenStudy (anonymous):

use the "FOIL" method multiply the first terms, then the 2 outer terms, then inner, then the last terms

OpenStudy (anonymous):

that takes forever tho is there a faster method

OpenStudy (anonymous):

Essentially, whenever you're solving quadratics or simplifying them, you have 3 rules. It sure beats simplifying every single term. a2-b2 = (a-b)(a+b) a2-2ab+b2 = (a-b)2 a2+2ab+b2 = (a+b)2 You can manually solve (x+8)(x+2) by splitting it up into x(x+2) + 8(x+2). That would give you x2 +2x +8x+16 which gives you x2+6x+16. ;

OpenStudy (anonymous):

still kinda confused

OpenStudy (anonymous):

do u mean a^2?